# Evaluating End-User Device Energy Models in Sustainability Reporting of Browser-Based Web Services

### v1

@@ 5.1. Model Accuracy and Error Growth (RQ1)
Figure 2. Measured versus modeled energy consumption for the DIMPACT framework. The red dashed line indicates perfect prediction ($y=x$).
Figure 3. Measured versus modeled energy consumption for the Digst framework. The stronger deviation from the reference line reflects systematic underestimation.
Figure 4. Relationship between task duration and absolute percentage error in DIMPACT energy estimates ($\beta=5.0\,\text{J s}^{-1},R^{2}=0.86$).
Figure 5. Relationship between task duration and absolute percentage error in Digst energy estimates ($\beta=11.8\,\text{J s}^{-1},R^{2}=0.96$).
5.1.1. Model Accuracy
Figure 2 compares measured energy with DIMPACT’s model estimates. The values are positively correlated, but the model systematically underestimates actual energy use for higher-energy tasks. Across all flows, the mean percentage error was $26.2\,\%$, the mean absolute percentage error (MAPE) was $28.7\,\%$, and the root-mean-square error (RMSE) was $25.5\,\text{J}$. The corresponding Digst comparison (Figure 3) revealed a stronger underestimation bias, with a mean error of $49.7\,\%$, MAPE of $49.7\,\%$, and RMSE of $38.1\,\text{J}$. While both models show moderate overall agreement with measured values, the deviation from the 1:11{:}1 reference line confirms that the constant-power assumption does not hold for realistic browser interactions.
5.1.2. Systematic Growth of Error with Duration.
To test whether model error accumulates with task length, we examined the relationship between task duration and absolute percentage error. For DIMPACT (Figure 4), a linear regression explained $86\,\%$ of the variance in absolute error ($R^{2}=0.86,F(1,165)=989.5,p<0.001$), with a slope of $5.0\pm 0.16\,\text{J s}^{-1}$. For Digst (Figure 5), the relationship was even stronger, with $96\,\%$ of the variance explained ($R^{2}=0.96,F(1,165)=3929,p<0.001$) and a slope of $11.8\pm 0.19\,\text{J s}^{-1}$. These results show that model deviations increase almost linearly with interaction length: longer user flows yield proportionally larger relative errors. The pattern demonstrates that small per-second inaccuracies in the $P_{\mathrm{avg}}\times t$ approximation accumulate over time, leading to systematic underestimation for extended browsing sessions. Together, these findings confirm that although Digst and DIMPACT capture the general order of magnitude of web-service energy use, neither provides accurate or duration-independent estimates under realistic usage conditions.
5.2. Factors Impacting Model Deviations (RQ2)
To determine which factors explain the deviations observed in RQ1, we incrementally extended the proportional $P\times t$ model with parameters representing hardware and web-service variability. All models were fitted without intercepts to preserve the physical relationship ($E_{\mathrm{model}}=P_{\mathrm{avg}}\times t$). Table 2 summarizes the progression in explanatory power.
5.2.1. Hardware and Category Effects.
The hardware-extended model ($E\sim t\!:\!C(\text{computerid})-1$) showed systematic differences in effective power between laptops ($R^{2}=0.960$), ranging from approximately 9 W to 12 W. These differences align with the measured distributions of average power across devices shown in Figure 6. Such hardware variation reflects both CPU generation and OEM design, and accounts for a substantial share of the total variance in energy use.
At the category level, user-flow differences explain an even larger portion of the remaining variance. Figure 7 illustrates that total energy consumption varies systematically between user flows, with news and navigation tasks exhibiting the highest medians. Corresponding differences in average device power (Figure 8) confirm that these effects stem from task intensity rather than duration alone. Together, these visual patterns substantiate the statistical improvements observed when adding category and hardware factors in the model-progression analysis (Table 2).
Figure 6. Distribution of average device power across laptops (C2–C5). Hardware-specific differences align with the coefficients in the duration-by-computer model ($R^{2}=0.960$), ranging from roughly 9 W for the most efficient to 12 W for the least efficient device.
@@ 5.2. Factors Impacting Model Deviations (RQ2)
Figure 8. Distribution of average device power across user flows. Power intensity varies by task type, confirming that flow-specific coefficients capture systematic rather than random variance.
5.2.2. Baseline and Hardware Effects
The global model ($E\!=\!10.15\times t$) achieved $R^{2}=0.946$, corresponding to an average effective power of 10.15 W. Introducing computer-specific coefficients increased the fit to $R^{2}=0.960$, revealing systematic power differences across devices (9.17–12.43 W). This reflects both CPU generation and OEM design variations.
5.2.3. Category and Interaction Effects.
Adding category-specific power factors improved the model to $R^{2}=0.973$, with typical fitted powers ranging from 8.4 W for navigation to 14.0 W for news tasks. Including detailed flow-level distinctions further increased $R^{2}$ to 0.978. When hardware and category were combined ($C(\text{category})\!:\!C(\text{laptop})$), the model explained $98.9\,\%$ of variance, and the full interaction model across flow and computer reached $R^{2}=0.994$. These results show that nearly all residual variance stems from identifiable and systematic differences in hardware and interaction type, rather than random error.
Table 1. Progression of OLS models explaining measured energy ($E$) from duration ($t$) and contextual factors (with intercept = 0).
@@ 5.2.2. Baseline and Hardware Effects
5.2.4. Interpretation.
The progression demonstrates that hardware and web-category parameters substantially increase both precision and accuracy of energy estimation. The resulting context-aware model nearly reproduces measured values ($R^{2}=0.994$), confirming that the fixed-power assumptions of current reporting frameworks cannot capture the heterogeneity of realistic web-service energy use.
6. Discussion
@@
The results show that deviations between modeled and measured energy are not random but scale systematically with interaction length, revealing a structural bias in the constant-power assumption of existing reporting models. The nearly perfect linear relation between Digst error and task duration confirms that this bias arises from the model formulation rather than from random measurement noise.
6.1. Improving Accuracy of Reporting Frameworks
The analysis in Section 5.2 showed that model deviations are systematic rather than random, driven primarily by reproducible differences in device power and service category. While incorporating hardware parameters can explain nearly all variance ($R^{2}=0.994$), this level of granularity is impractical for sustainability reporting. A more feasible improvement lies in introducing a limited set of representative service categories —such as shopping, booking, navigation, and news— which capture much of the same systematic variation ($R^{2}=0.973$). This categorical differentiation could substantially increase model precision while preserving the simplicity and comparability that reporting frameworks require. Such category grouping could therefore enhance the transparency and fairness of sustainability reporting without introducing additional measurement burden or model complexity.
6.2. Toward Reproducible and Transparent Sustainability Assessment
Digital sustainability models should incorporate reproducibility standards and sensitivity analyses that explicitly address hardware and user variance. Transparent reporting of assumptions, parameter sources, and version-controlled model dependencies would enable consistent comparisons across services and over time. The open repository accompanying this paper provides all datasets and scripts necessary to reproduce the results, including detailed software version information to ensure full experimental repeatability.
6.3. End-User Devices
@@
Our measured laptop powers (9–13 W) align closely with modeled estimates in Malmodin’s long-term household measurements (malmodin2023, 19), which report average powers between 6 W and 12 W with short-term fluctuations from 3 W to 25 W. In addition, Malmodin’s results align well with the 10 W average derived in our best-fit model, which is realistic for browser-based workloads and representative of typical laptop energy behavior.
@@
In 2013, Schien et al. (Schien2013, 27) observed average power dissipation of a laptop to be 32–37 W [^3] In the supplementary material, the energy consumption for 10 minutes of reading text and video watching was reported as 19.15 kJ and 22.01 kJ, respectively., and they found no statistically significant variation from idle power when browsing text. This does not align with our results. For comparison, see Table 2, which lists the Energy Star measured idle power $P_{idle}$ (hp_elitebook_830g6_2020, 17, 5, 6, 4), the minimum observed power during execution, and the best-fit power average. The best-fit power average is significantly higher than the Energy Star idle power and, in some cases, twice as high.
@@
Table 2. For each laptop, the Energy Star $P_{idle}$, minimum observed power during execution, and the best-fit power dissipation average across all user flows.
No. Laptop $\mathbf{P}_{\mathbf{Idle}}$ Min Best-fit P
@@
Malmodin also noted that Apple’s M-series architectures achieve averages below 6 W, and subsequent silicon generations, such as the M4 and newer Intel hybrid designs, are expected to further reduce typical laptop power dissipation. Thus, as hardware efficiency continues to improve, periodic recalibration of model parameters will be necessary to ensure that reported end-user power values remain representative of actual device behavior.
While our measured laptop powers (9–13 W) correspond well to these empirically grounded estimates, they are markedly lower than the fixed constants used in current sustainability reporting frameworks. For instance, the Danish Agency of Digital Government’s Digst model assumes 22 W, and the DIMPACT framework applies 15 W for browser-based scenarios. These simplified constants serve comparability purposes but appear conservative relative to modern laptop behavior, leading to systematic overestimation of end-user energy use in typical web interactions. Because both under- and overestimation of average power compound proportionally with interaction duration, any mismatch between assumed and actual device power introduces systematic rather than random error in total energy estimates. This observation suggests a need to explore alternative formulations that decouple energy estimation from pure duration, for instance by distinguishing a time-dependent baseline component from an activity-dependent component that captures variations in user interaction intensity.
6.4. Broader Context
@@
Beyond laptop-based web interactions, similar mechanisms shape the energy profiles of other end-user workloads such as streaming and video communication. While this study focused on browser-based interactions, similar principles apply to other classes of end-user workloads, including streaming and video communication. Herglotz et al. (Herglotz2025, 16) demonstrated that energy measurement of video workloads requires precise synchronization between playback and power sampling. Our observations for the TV2 flow, which contained embedded autoplay videos, show markedly higher average power and total energy consumption than non-video pages, consistent with the controlled findings of Herglotz et al. (Herglotz2025, 16). Whereas their work isolated and synchronized video playback under laboratory conditions, our setup captured realistic end-user browsing sessions in which short videos appeared as part of ordinary news-site activity. Together, these perspectives confirm that even brief embedded videos can dominate device-level energy use, highlighting the need for category-specific modeling in sustainability reporting.
7. Limitations and Validity Considerations
To evaluate the robustness and scope of our findings, we examined potential threats to validity related to model structure, participant behavior, and experimental coverage.
Figure 9. Average measured energy across repeated runs (1–3). Energy decreased from Run 1 to Run 2 and remained stable thereafter, indicating a short learning effect without fatigue. Error bars represent standard deviations across participants.
Table 3. Number of distinct participants per flow–computer combination, showing balanced coverage across conditions.
Flow ID C2 C3 C4 C5
Aa 6 5 4 4
Ab 7 5 5 5
Ba 6 5 5 5
Bb 7 4 4 5
Ca 5 5 5 5
Cb 6 5 5 5
Da 6 5 5 5
Db 6 5 5 5
7.1. Model Structure and Bias
The proportional $P\times t$ formulation used in simplified reporting models inherently amplifies any misestimation of average device power. If the assumed power value is too high or too low, the resulting energy estimates accumulate a systematic bias proportional to task duration, as observed in Section 5.1. Consequently, even small parameter errors can produce large deviations for long-running interactions, limiting the external validity of constant-power approximations.
7.2. Participant Learning and Fatigue
Repeated participation may influence execution efficiency and thereby measured energy use. To mitigate such order effects, both the sequence of user flows and the allocation of laptops were randomized under constraints that ensured full coverage. Average energy use across all userflows declined from the first run (1453 J) to the second run (905J ) and then stabilized at (836 J) in the third run, see Figure 9 for the per userflow energy consumption. This pattern indicates a brief learning phase but no subsequent fatigue or performance drift. Because the learning effect was consistent across participants and evenly distributed through randomization, its influence on model validity is considered minimal.
7.3. Coverage and Inter-User Variance.
The experiment crossed eight user flows with four laptop platforms, resulting in a balanced dataset in which each flow–computer combination was represented by at least four and up to seven distinct participants (Table 3). This design ensured that device- and category-level effects were not driven by single-user behavior. Analysis of participant-level random effects in the mixed model (Section 5.2) yielded an intraclass correlation coefficient (ICC) of 0.012, indicating that only 1.2 % of total variance originated from between-participant differences. Thus, inter-user variability was negligible, and statistical power was sufficient for stable estimation of category- and device-level effects.
7.4. External Validity
The study focused on browser-based, non-streaming web services executed on four laptop platforms representative of common Intel generations and OEM designs. This scope ensured comparability across devices and categories but limits the generality of the findings in three respects. First, the tested service categories—shopping, booking, navigation, and news—represent only a subset of web interactions; extending the analysis to other categories such as streaming, productivity, or AI-assisted services would clarify the broader applicability of the results. Second, the dataset covers four laptop models, providing internal consistency but limited diversity across architectures. Expanding the hardware base to include a wider range of contemporary Energy Star–rated laptops would strengthen the trustworthiness of the derived power parameters. Third, the focus on laptops excludes other end-user devices such as mobile phones, tablets, and high-performance desktops, which may differ in baseline power and dynamic scaling behavior. All interactions were performed on live websites under real network conditions, ensuring ecological validity for laptop-based browsing. Consequently, the results are representative of realistic end-user conditions within this scope, while broader generalization to additional service types, hardware generations, and device classes should be examined in future studies.
8. Conclusion & Future Work
@@
This study show that the constant-power assumptions used in digital sustainability reporting frameworks such as Digst and DIMPACT may systematically overestimate end-user device energy use. Empirical measurements of realistic browser-based interactions revealed typical laptop powers between 9 W and 13 W—substantially lower than the 22 W and 15 W values assumed in these models. As a result, model bias increases proportionally with task duration, producing structural rather than random error.
@@
Assuming a near-perfect constant-power exists, we found that intorducing hardware and service-category parameters increases model precision markedly. This showed that most variance originates from reproducible, context-dependent differences in devices and workloads. These findings suggest that sustainability reporting can remain simple yet become substantially more accurate by adopting, e.g., category-specific power parameters.
All data, analysis scripts, and software versions are openly provided, enabling full reproducibility and future extensions of this work. Together, these results establish a methodological foundation for transparent, empirically grounded models of end-user device energy use, supporting more credible and comparable digital sustainability reporting.
Future work should aim to enhance the accuracy, transparency, and representativeness of end-user device energy models. First, expanding the hardware base to include a broader selection of contemporary laptops would improve the trustworthiness of derived parameters. Second, building on the evidence that constant-power assumptions systematically overestimate device energy, new models should explore ways to reduce systematic error by relating energy use to task context, interaction type, and hardware characteristics.
Although this study focused on browser-based interactions, it may be beneficial to explore whether the same principles could be extended to other software categories—such as streaming, productivity, and AI-assisted applications—to examine whether similar proportional relationships hold. Further research is required to refine category definitions and develop standardized, fair user flows to support reproducible end-user device energy assessments.

### v2

@@ IV-A Model Accuracy and Error Growth (RQ1)
Figure 3: Measured (x-axis) versus Digst’s modeled energy consumption (y-axis). The red dashed line indicates perfect prediction ($y=x$). All points lie above the perfect prediction; Digst overestimates the energy consumption.
@@
IV-A1 Model Accuracy
Figure 2 compares measured energy with DIMPACT’s model estimates. The values are positively correlated, but the model systematically overestimates actual energy use for higher-energy tasks. Across all flows, the mean percentage error was $26\,\%$, the mean absolute percentage error (MAPE) was $28.55\,\%$, and the root-mean-square error (RMSE) was $653\,\text{J}$.
The corresponding Digst comparison (Figure 3) revealed a stronger overestimation bias: all observations lie above the perfect-prediction line, yielding mean percentage error of $49.54\,\%$, MAPE of $49.54\,\%$, and RMSE of $1457\,\text{J}$.
While both models show moderate agreement with measured values, the deviation from the perfect prediction line confirms that the constant-power assumption does not hold for realistic browser interactions. The $R^{2}$-values are negative (Digst: 5.6-5.6, DIMPACT: 0.3-0.3) meaning that these predictions perform worse than always using the mean measured energy.
Figure 5: Absolute error in predicted energy (y-axes) versus task duration (x-axes) for Digst (blue) and DIMPACT (yellow). In both cases, the fitted through-origin model shows duration-proportional overestimation, with slopes $5.0\,\text{J/s}$ for DIMPACT and $11.8\,\text{J/s}$ for Digst.
IV-A2 Systematic Growth of Error with Duration.
To assess whether error accumulates with task length, we regressed absolute energy error on task duration (Figure 5). DIMPACT (yellow markers and dashed line) shows a strong duration effect ($R^{2}=0.86$, $F(1,164)=983.5$, $p<0.001$) with slope $\beta=5.0\pm 0.16\,\text{J/s}$, while Digst (blue markers and dashed line) is stronger ($R^{2}=0.96$, $F(1,164)=3905$, $p<0.001$) with $\beta=11.8\pm 0.19\,\text{J/s}$. In conclusion, deviations increase approximately linearly with interaction duration; per-second miscalibrations in $P_{\mathrm{avg}}!\times t$ therefore compound into systematic overestimation for longer browsing sessions.
IV-B Factor Effects on Model Deviations (RQ2)
To determine which factors explain the deviations observed in RQ1, we incrementally extended the constant power model with parameters representing hardware and web-service variability. All models were fitted with intercepts in (0,0) to reflect the nature of energy usage for the formula $E=t\times P$: zero energy is used in tasks of zero seconds duration. Table I summarizes the progression in explanatory power, were a (0,0) intercept is enforced by a ”-1” in the formula and ”C(criteria)” indicates the partitioning criteria for the data set before training.
TABLE I: Progression of OLS models explaining measured energy ($E$) from duration ($t$) and contextual factors (with intercept = 0).
@@ IV-B Factor Effects on Model Deviations (RQ2)
IV-B1 Baseline
The global model ($E\!=\!10.15\times t$) achieved $R^{2}=0.946$, corresponding to an average power of 10.15 W.
IV-B2 Individual Effects
Introducing computer-specific coefficients increased the fit to $R^{2}=0.960$, revealing systematic power-model differences across devices (9.17–12.43 W). These differences align with the measured power across devices shown in Figure 6. Such hardware variation reflects both CPU generation and OEM design, and accounts for a substantial share of the total variance in energy use. Adding category-specific power factors improved the model to $R^{2}=0.973$, with typical fitted powers ranging from 8.5 W for navigation to 14 W for news tasks. Using flow-level distinction further increased $R^{2}$ to 0.978. Corresponding differences in average device power (Figure 7) confirm that these effects stem from task intensity rather than duration alone.
Figure 6: Distribution of average device power across laptops (L1–L4). Hardware-specific differences align with the coefficients in the duration-by-computer model ($R^{2}=0.960$), ranging from roughly 9 W for the most efficient to 12 W for the least efficient device.
Figure 7: Distribution of average device power across user flows. Power intensity varies by task type, confirming that flow-specific coefficients capture systematic rather than random variance.
IV-B3 Interaction Effects.
When hardware and category were combined ($C(\text{category})\!:\!C(\text{laptop})$) obtained $R^{2}=0.989$. and the full interaction model across flow and computer reached $R^{2}=0.994$. These results show that nearly all residual variance stems from identifiable and differences in hardware and interaction type, rather than random error.
V Threats to Validity
We discuss threats to validity related to participant behavior and execution order, the operationalization of services and flows, and the generalizability of the evaluated workload and hardware scope.
V-A Internal validity: order effects and participant behavior
Repeated participation may change execution efficiency and thereby measured energy consumption. To mitigate order effects, we randomized both the sequence of user flows and the assignment of laptops under constraints that ensured coverage across conditions. Average energy decreased from Run 1 (1453 J) to Run 2 (905 J) and then stabilized in Run 3 (836 J); Figure 8 shows the corresponding pattern at the flow level. This indicates a brief learning effect but no evidence of fatigue or performance drift. Each flow–computer combination was executed by multiple participants (min. 4, max. 7), and between-participant variance in the mixed-effects analysis was small (ICC =0.012=0.012), suggesting that idiosyncratic user behavior had limited influence on the results.
Repeated participation can affect execution efficiency and measured energy. We have mitigated order effects by randomizing both flow order and laptop assignment while ensuring coverage across conditions. Mean energy dropped from Run 1 (1453 J) to Run 2 (905 J) and then stabilized in Run 3 (836 J) (Figure 8), indicating a brief learning effect but no fatigue or drift. Each flow–computer condition included multiple participants (min. 4, max. 7), and participant-level variance was small (ICC =0.012=0.012), indicating that individual differences had little impact on the results.
V-B Construct validity: representing a service
Each service was evaluated using one essential user flow. This is appropriate for controlled model comparison, but service-level reporting may require aggregating multiple essential flows to represent typical use.
V-C External validity: workload, hardware, and device scope
The study evaluates browser-based, non-streaming services on four laptop platforms spanning common Intel generations and OEM designs. Generalization is limited in three respects: (i) the evaluated categories (shopping, booking, navigation, news) represent only a subset of web interactions; (ii) the dataset includes four laptop models, limiting vendor and architectural diversity; and (iii) results are specific to laptops and may not transfer to phones, tablets, or high-performance desktops with different baseline power and scaling behavior. All interactions were executed on live websites under real network conditions, supporting ecological validity for laptop-based browsing within this scope.
Figure 8: Average measured energy across repeated runs (1–3). Energy decreased from Run 1 to Run 2 and remained stable thereafter, indicating a short learning effect without fatigue. Error bars represent standard deviations across participants.
VI Discussion
@@
The results show that deviations between modeled and measured energy are not random but scale systematically with interaction length, revealing a systematic bias in the constant-power assumption of existing reporting models. The nearly perfect linear relation between Digst error and task duration confirms that this bias arises from the model formulation rather than from random measurement noise.
VI-A Power Average of modern End-User Devices
Our analysis of factors effect (Section IV-B) started from the point of creating a new best-fit model based on the available data. This increased $R^{2}$ from a negative number, i.e., using the average measured energy usage is a better choice, to $94.6\%$. This shows that –when using an energy model with constant power assumption– it in essential to calibrate the power average. Therefore, it is essential to compare our measured laptop power usage with that of the existing literature.
The measured average laptop powers in our study (9–13 W) are substantially lower than the fixed constants currently used in sustainability reporting frameworks. For example, the Danish Agency for Digital Government (Digst) assumes 22 W, and DIMPACT applies 15 W for browser-based scenarios. Consequently, using inaccurate power constants can introduce duration-scaled bias rather than random estimation noise.
@@ VI-A Power Average of modern End-User Devices
Our observations align closely with those in Malmodin’s recent study on long-term household measurements [25], which report average powers between 6 W and 12 W with short-term fluctuations from 3 W to 25 W. In addition, Malmodin’s results align well with the 10 W average derived in our best-fit model, which is realistic for browser-based workloads and representative of typical laptop energy behavior.
@@
In 2013, Schien et al. [6] observed average power dissipation of a laptop to be 32–37 W [^7] Supplementary material [6] report energy consumption for 10 minutes of reading text and video watching as 19.15 kJ and 22.01 kJ, respectively., and they found no statistically significant variation from idle power when browsing text. This does not align with our results. For comparison, see Table II, which lists the Energy Star measured idle power $P_{idle}$ [26, 27, 28, 29], the minimum observed power during execution, and the best-fit power average. Our best-fit power average is significantly higher than the Energy Star idle power and, in some cases, twice as high.
@@
TABLE II: The laptops Energy Star $P_{idle}$, minimum observed power during execution, and the best-fit power average.
No. Laptop $\mathbf{P}_{\mathbf{Idle}}$ Min. $P$ Best-fit $P$
@@
Malmodin also noted that Apple’s M-series architectures achieve averages below 6 W, and subsequent silicon generations, such as the M4 and newer Intel hybrid designs, are expected to further reduce typical laptop power dissipation.
Overall, our measurements are consistent with ongoing hardware efficiency improvements, implying that reporting models should be periodically recalibrated to keep assumed end-user power values representative of contemporary devices. Alternatively, this suggests a need to explore alternative formulations that decouple energy estimation from pure duration, for instance by distinguishing a time-dependent baseline component from an activity-dependent component that captures variations in user interaction intensity.
VI-B Improving Accuracy of Reporting Frameworks
The purpose of reporting models is to enable fair and practically applicable comparisons of services. Our analysis of factor effects (Section IV-B) started by estimating a new best-fit model from the available data. This increased $R^{2}$ from a negative value—indicating that the mean measured energy consumption would provide a better predictor—to $94.6\%$. Under the assumption that systematic error can be controlled, model precision becomes a central concern.
Across the evaluated alternatives, the model combining hardware characteristics with individual user flows achieved the highest explanatory power ($R^{2}=99.4\%$). In contrast, the scientific basis underlying the national models, as discussed by Preist et al., proposes averaging across Energy Star measurement results (i.e., across certified “green” devices) to reflect end-user device heterogeneity [4]. This emphasis on device diversity is compatible with the objective of practical reporting. However, operationalizing it in our setting would require measuring the same user flows on a sufficiently broad set of labelled devices to capture the plausible diversity of end-user hardware. Such a measurement effort may be infeasible, and methods for reducing this burden are beyond the scope of the present paper. Future work should therefore investigate the extent of device sampling required for robust reporting, and how this sampling can be achieved cost-effectively.
In some contexts, direct energy measurement may be impractical (e.g., for small services with limited user time). In our data, categories alone increased $R^{2}$ to $97.3\%$, suggesting that refining DIMPACT’s service categories could support model-based estimation when measurement is not feasible. While each category here included only two services, the results indicate that category-based refinement can improve model fit and support fairer comparisons among functionally similar services, akin to EU energy labelling.
VI-C Video Streaming
@@
Although this study focuses on browser-based interactions, similar effects are expected in other end-user workloads, including streaming and video communication. In the TV2 flow, which included embedded autoplay video, we observed higher average power than on non-video pages, consistent with the controlled results of Herglotz et al. [23]. Whereas their study isolated and synchronized video playback under laboratory conditions, our setup captures realistic browsing sessions. Taken together, the results indicate that even brief embedded videos can dominate device-level energy use, motivating category-specific modeling in sustainability reporting.
VII Conclusion & Future Work
@@
This study demonstrates that the constant-power assumptions used in digital sustainability reporting frameworks (e.g., Digst and DIMPACT) can systematically overestimate end-user device energy use. Across realistic browser-based interactions, we measured typical laptop powers of 9–13 W, substantially below the 22 W and 15 W values assumed by these models. Because energy is estimated as $P\times t$, any power mismatch induces a duration-proportional bias, yielding systematic rather than random error.
@@
Using a constant-power baseline, adding hardware and service-category parameters substantially improved model precision, showing that much of the variance reflects reproducible, context-dependent differences in devices and workloads. Thus, reporting can remain lightweight while becoming more accurate by adopting calibrated category-specific (and, where feasible, device-specific) power parameters.
Future work would benefit from pursuing three directions. First, expand the hardware base to a larger and more diverse set of contemporary laptops to improve the representativeness of derived parameters. Second, reduce systematic overestimation by moving beyond global constants and modeling energy as a function of task context, interaction type, and hardware characteristics. Third, evaluate the approach on additional software categories (e.g., streaming, productivity, and AI-assisted applications) and develop refined category definitions and standardized, fair user flows to support reproducible assessment.